# Robustness of dark energy phenomenology across different parameterizations

### v2

@@ 3.1 Thawing quintessence with minimal coupling
$S=\int d^{4}x\sqrt{-g}\left[\frac{1}{2}M_{\mathrm{\rm Pl}}^{2}R-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\varphi\partial_{\nu}\varphi-V(\varphi)\right]+S_{\rm m}.$ (3.1)
$M_{\rm Pl}$ is the reduced Planck mass, $g$ is the determinant of the metric $g_{\mu\nu}$, $R$ is the Ricci curvature scalar, $V(\varphi)$ is the scalar field potential, and $S_{\rm m}$ is the action for matter.
@@
$H^{2}\equiv\left(\frac{\dot{a}}{a}\right)^{2}=\frac{1}{3M^{2}_{\rm Pl}}\left[\frac{1}{2}\dot{\varphi}^{2}+V(\varphi)+\rho_{\rm m}\right],$ (3.4)
@@
where throughout this paper, we will give $m^{2}$ in units of $(H_{0}/h)^{2}$ and $V_{0}$ in units of $M^{2}_{\rm Pl}(H_{0}/h)^{2}$. This model has been implemented in hi_class [118, 119], where the initial conditions are chosen to generically set $\dot{\varphi}_{ini}=0$ and $\varphi_{ini}$ is tuned to satisfy the equations of motion while the other parameters are varied (see [12] for an analytic expression for $\varphi_{ini}$ in terms of model and cosmological parameters).
This model can capture the entire phenomenology of thawing quintessence because, when we have a positive mass term in Eq. (3.11), it behaves as “slow-roll” thawing quintessence [120], which has been extensively studied and showed to converge to a universal phenomenological behavior given by an approximately linear evolution in the equation of state. This behavior is well-described by $dw/da|_{a=1}\approx 1.5[1+w(a=1)]$ (e.g. [120, 121, 112, 82, 122, 123]). When the mass is negative, this corresponds to “hilltop” quintessence [124, 125, 84, 82, 75], which has much more varied dynamical behavior. If $\left|m^{2}\right|$ is small, the resulting $w(a)$ approximates the linear behavior of the positive branch. On the other hand, if $\left|m^{2}\right|$ is large, $w(a)\simeq-1$ for much of cosmic history, but can evolve very rapidly and non-linearly at more recent times depending on the choice of model parameters, allowing it to “sweep” over the $(w_{0},w_{a})$ parameter space. In addition to the above arguments, we can also understand why this model is able to subsume the phenomenology of thawing quintessence in the following way. As any arbitrary analytic potential can be expand in a Taylor series (e.g. [124, 82, 125]), the potential given in Eq. (3.11) can be seen as the leading order terms for a huge variety of scalar field potentials that broadly fall under either the “slow-roll” and “hilltop” types; and in the regime of field space for which quintessence can play the role of cosmological dark energy, these are the terms that will be relevant.
@@ 3.2 Thawing quintessence with non-minimal coupling
where $\xi$ is the dimensionless coupling coefficient between the scalar field $\varphi$ and gravity $R$. This introduces a modification into the gravitational sector (while keeping the matter sector of minimally coupled quintessence the same) [^2] Other kinds of non-minimal couplings such as a direct coupling between the scalar field and dark matter are also possible (e.g. [126, 127, 128, 129, 76]), but here we will restrict our focus only to non-minimal couplings with respect to gravity., which changes the corresponding scalar field equation of motion as well as the standard Friedmann equation from GR:
@@
$H^{2}=\frac{1}{3M_{\text{\rm Pl}}^{2}\left(1-\xi\varphi^{2}\right)}\left[\frac{1}{2}\dot{\varphi}^{2}+V(\varphi)+3H\xi\varphi\dot{\varphi}+\rho_{m}\right].$ (3.8)
@@
can have a recent phantom crossing along with very sharp dynamical evolution, and they display a significant improvement to the fit of the data over $\Lambda$CDM or quintessence [13].
@@
where, as before, each of these terms can be thoughts of as the leading order terms in a Taylor expansion. This model has been implemented in hi_class [118, 119], where the initial conditions are chosen such that $\varphi_{ini}=0$ and $\dot{\varphi}_{ini}\simeq 0$. We tune $V_{0}$ to satisfy the equations of motion, finding that $V_{0}\simeq 3H_{0}^{2}\Omega_{\varphi}$ is a robust starting point for the numerical solvers. Similar to before, the units of $V_{0}$ are given in $M^{2}_{\rm Pl}(H_{0}/h)^{2}$, the units of $\beta$ are given in $M_{\rm Pl}(H_{0}/h)^{2}$, and the units of $m^{2}$ are given in $(H_{0}/h)^{2}$. Of course, as a modified gravity model, this model can suffer from instabilities (see e.g. [130, 131]). In this paper we will be primarily concerned with predictions for background quantities, but we use hi_class to perform stability checks and ensure viability of all the parameter choices considered here by computing linear cosmological perturbations for each model.
In analogy with thawing quintessence above, we can refer to this as Non-Minimal Thawing Quintessence as it will phenomenologically capture the behavior of dark energy models with this kind of non-minimal coupling and an arbitrary potential. For example, it captures the behavior of non-minimal models with an exponential potential and with a hilltop potential which have been studied in [13, 14, 47], as again this form can be seen as corresponding to a Taylor expansion in the potential. In [13], it was demonstrated that such models can saturate much of the thawing region of the CPL $(w_{0},w_{a})$ plane. See also [20] for a full Bayesian analysis where it is found that this type of non-minimally coupled dark energy model is strongly favoured over $\Lambda$CDM. However, in this paper, we will focus on the phenomenology and whether or not the aforementioned dark energy parametrizations are equipped to capture it.
@@ 5.2 Reproducing dark energy phenomenology
As can be seen in Fig. (3), the fitting procedure we have used reproduces the predicted cosmological observables of both thawing quintessence and non-minimal thawing quintessence very accurately for all parameterizations considered. While there are some differences between the parameterizations, given that we have $\text{DoF}\simeq 20$ here, when examining the $\chi^{2}$ determined from minimizing Eq. (3.12) we consistently have that $\chi^{2}/\text{DoF}<1$ or $\simeq 1$ across the board regardless of the the dark energy model or parameterization considered. For both the positive branch of thawing quintessence and the non-minimal thawing quintessence model, $\sim 99.99\%$ of the models for all parameterizations have $\chi^{2}/\text{DoF}<1$. For the hilltop branch it is a bit more interesting as the JBP parameterization is such that $\sim 99\%$ of the models have $\chi^{2}/\text{DoF}<1$, whereas for the other parameterizations it is $\sim 95\%$ of the models that have $\chi^{2}/\text{DoF}<1$.
@@ 5.3 Data constraints and viability of the dark energy models
This is confirmed when we compare the various parametrizations and thawing quintessence in terms of their ability to fit the complete (uncompressed) data. To do so, we find the best fitting model of thawing quintessence (i.e. that maximizes the likelihood ${\cal L}$ or, alternatively, minimizes $\chi^{2}=-2\ln{\hat{\cal L}}$ for the given datasets [^5] Here, we use the lite version of Planck likelihood to do the minimization as it has substantially fewer parameters than the full likelihood and is thus more efficient.). We do the same with $\Lambda$CDM and with the CPL, JBP, EXP, and BA parameterizations. The best-fit $\chi^{2}$ can can be seen in Table (2) and show that both thawing quintessence and all parametrization can fit the data better than $\Lambda$CDM, in agreement with [29]. In particular, while the scalar field improves the $\chi^{2}$ of $\Lambda$CDM only by $\Delta\chi^{2}=-2.9$, the parametrizations obtain $\Delta\chi^{2}=-6.7,\,-6.9,\,-7.1$, for CPL, EXP and BA. The worst performing parametrization is JBP, with $\Delta\chi^{2}=-4.9$, and this worse fit is the reason why in Fig. (4), for JBP, the wedge of the theory projection has a bigger overlap with the data constraints: both JBP and thawing quintessence are relatively close in terms of their ability to fit the expansion history data, meaning that when thawing quintessence is projected into JBP parameter space there will be more overlap than with the other parameterizations which are further away from thawing quintessence in terms of their ability to fit the data.
@@
There are, however, a couple of points worth briefly commenting on. Naively, if we use the Akaike Information Criterion (AIC), which is given by ${\rm AIC}=2k+\chi^{2}$, where $k$ is the number of parameters in the model, we find that $\Delta{\rm AIC}\simeq 12.5$ because the non-minimal model has 5 parameters (the 4 mentioned above plus one additional parameter $\dot{\phi}_{ini}$ for the initial conditions) while $\Lambda$CDM has only one ($\Lambda$). Even accounting for the additional number of parameters, this is still a notable improvement that, as mentioned above, merits further investigation.
Another point to consider is that, while Section (5.2) demonstrated that the various dark energy parameterizations can all accurately reproduce the phenomenology of both minimal and non-minimal thawing quintessence, the fact that the non-minimal model can achieve a lower $\chi^{2}$ than any of the parameterizations does show that there is some information that is being lost when the full dark energy model is projected into the parameterizations’ parameter spaces. However, this makes sense as fitting directly to the observables using Eq. (3.12) was only fit to background observables that represent compressed versions of the full data (with the exception of the BAO data which requires no compression). The biggest difference between the compressed (background) observables and the full data will be in the CMB power spectra and CMB lensing data from Planck and ACT, where one must compute the full linear cosmological perturbations to compare with the data. And indeed, while we find that the non-minimal model reproduces similar $\chi^{2}$ with respect to the SNe and BAO components of the full data when compared with the various parameterizations, the improvement in overall $\chi^{2}$ ($\Delta\chi^{2}\simeq 4$ in the case of the DES-Y5 dataset) over those of the various parameterizations was largely driven by the the Planck and ACT components of the data. So while these parameterizations clearly perform exceptionally well at reproducing dark energy phenomenology at the background level, it is perhaps not surprising that there is some information lost at the perturbative level.
@@ 6 Discussion
Our analysis reveals that assessing the viability of these dynamical dark energy models is very robust to the choice of parameterization. By mapping the dynamical dark energy models into the $(w_{0},w_{a})$ parameter space for each parameterization, we demonstrated both that our conclusions regarding their viability (or how well they describe the data) are not overly subject to particular choices of parameterization and also that all of these various parameterizations effectively capture the phenomenology of the full dark energy models. This demonstrates that, at least for these very broad and simple kinds of microphysical dark energy models considered here, any of these choices of parameterization are sufficient to interpret the data and make assessments regarding the viability of such dark energy proposals.
@@ Appendix A Compressed data for theory priors
This is exactly analogous to what is done in the study of inflation models. We make observations that constrain the tensor-to-scalar ratio $r$ and the scalar spectral index $n_{s}$, and then use these constraints to determine the observationaly favored region of the $(r,n_{s})$ plane. To evaluate whether or not specific inflationary models are viable, we compute their predictions for $(r,n_{s})$ and compare these to those constraints (e.g. [147, 148, 149, 150, 151]). However, unlike with inflation, as we have seen there are a number of distinct parameterizations for dark energy and these parameters are a bit further from the actual observables. So, in order to actually determine how well a dark energy model maps onto these constraints, we must determine what a dark energy model would predict for $(w_{0},w_{a})$ by using a procedure that mimics what was done for determining where the data itself lies in terms of the parameters. In the dark energy literature, it is common to generate a dark energy model’s equation of state and “fit” this equation of state over some recent range of redshifts using one of the parameterizations directly as a proxy for determining where a model might lie on the $(w_{0},w_{a})$ plane. However, as discussed in [82, 13, 12], because we do not measure the equation of state directly, doing this can often lead to highly misleading results. In other words, we must generate the actual observables for a given dark energy model, and then determine what $(w_{0},w_{a})$ parameters best reproduce those observables given the the known errors and uncertainties of the data. This will allow us to make an “apples to apples” comparison.

### v3

@@ 3.1 Thawing quintessence with minimal coupling
$S=\int d^{4}x\sqrt{-g}\left[\frac{1}{2}M_{\mathrm{\rm P}}^{2}R-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\varphi\partial_{\nu}\varphi-V(\varphi)\right]+S_{\rm m}.$ (3.1)
$M_{\rm P}$ is the reduced Planck mass, $g$ is the determinant of the metric $g_{\mu\nu}$, $R$ is the Ricci curvature scalar, $V(\varphi)$ is the scalar field potential, and $S_{\rm m}$ is the action for matter.
@@
$H^{2}\equiv\left(\frac{\dot{a}}{a}\right)^{2}=\frac{1}{3M^{2}_{\rm P}}\left[\frac{1}{2}\dot{\varphi}^{2}+V(\varphi)+\rho_{\rm m}\right],$ (3.4)
@@
where throughout this paper, we will give $m^{2}$ in units of $(H_{0}/h)^{2}$ and $V_{0}$ in units of $M^{2}_{\rm P}(H_{0}/h)^{2}$. This model has been implemented in hi_class [118, 119], where the initial conditions are chosen to generically set $\dot{\varphi}_{ini}=0$ and $\varphi_{ini}$ is tuned to satisfy the equations of motion while the other parameters are varied (see [12] for an analytic expression for $\varphi_{ini}$ in terms of model and cosmological parameters).
This model can capture the entire phenomenology of thawing quintessence because, when we have a positive mass term in Eq. (3.5), it behaves as “slow-roll” thawing quintessence [120], which has been extensively studied and showed to converge to a universal phenomenological behavior given by an approximately linear evolution in the equation of state. This behavior is well-described by $dw/da|_{a=1}\approx 1.5[1+w(a=1)]$ (e.g. [120, 121, 112, 82, 122, 123]). When the mass is negative, this corresponds to “hilltop” quintessence [124, 125, 84, 82, 75], which has much more varied dynamical behavior. If $\left|-m^{2}\right|$ is small, the resulting $w(a)$ approximates the linear behavior of the positive branch. On the other hand, if $\left|-m^{2}\right|$ is large, $w(a)\simeq-1$ for much of cosmic history, but can evolve very rapidly and non-linearly at more recent times depending on the choice of model parameters, allowing it to “sweep” over the $(w_{0},w_{a})$ parameter space. In addition to the above arguments, we can also understand why this model is able to subsume the phenomenology of thawing quintessence in the following way. As any arbitrary analytic potential can be expand in a Taylor series (e.g. [124, 82, 125]), the potential given in Eq. (3.5) can be seen as the leading order terms for a huge variety of scalar field potentials that broadly fall under either the “slow-roll” and “hilltop” types; and in the regime of field space for which quintessence can play the role of cosmological dark energy, these are the terms that will be relevant.
@@ 3.2 Thawing quintessence with non-minimal coupling
where $\xi$ is the dimensionless coupling coefficient between the scalar field $\varphi$ and gravity $R$. This introduces a modification into the gravitational sector (while keeping the matter sector of minimally coupled quintessence the same) [^2] Other kinds of non-minimal couplings such as a direct coupling between the scalar field and dark matter are also possible (e.g. [126, 127, 128, 129, 76, 130]). , which changes the corresponding scalar field equation of motion as well as the standard Friedmann equation from GR:
@@
$H^{2}=\frac{1}{3M_{\text{\rm P}}^{2}\left(1-\xi\varphi^{2}\right)}\left[\frac{1}{2}\dot{\varphi}^{2}+V(\varphi)+3H\xi\varphi\dot{\varphi}+\rho_{m}\right].$ (3.8)
@@
as was done in [13] can have a recent phantom crossing along with rapid dynamical evolution and display a significant improvement to the fit of the data over $\Lambda$CDM or quintessence.
@@
where, as before, each of these terms can be thoughts of as the leading order terms in a Taylor expansion and the coefficients can be positive or negative. This model has been implemented in hi_class [118, 119], where the initial conditions are chosen such that $\varphi_{ini}=0$ and $\dot{\varphi}_{ini}\simeq 0$. We tune $V_{0}$ to satisfy the equations of motion, finding that $V_{0}\simeq 3H_{0}^{2}\Omega_{\varphi}$ is a robust starting point for the numerical solvers. Similar to before, the units of $V_{0}$ are given in $M^{2}_{\rm P}(H_{0}/h)^{2}$, the units of $\beta$ are given in $M_{\rm P}(H_{0}/h)^{2}$, and the units of $m^{2}$ are given in $(H_{0}/h)^{2}$. Of course, as a modified gravity model, this model can suffer from instabilities (see e.g. [131, 132]). In this paper we will be primarily concerned with predictions for background quantities, but we use hi_class to perform stability checks and ensure viability of all the parameter choices considered here by computing linear cosmological perturbations for each model.
In analogy with thawing quintessence above, we can refer to this as Non-Minimal Thawing Quintessence as it will phenomenologically capture the behavior of dark energy models with this kind of non-minimal coupling and an arbitrary potential. For example, it captures the behavior of non-minimal models with a simple exponential potential and with a hilltop potential which have been studied in [13, 14, 47], as again this form can be seen as corresponding to a general Taylor expansion in the potential. In [13], it was demonstrated that such models can saturate much of the thawing region of the CPL $(w_{0},w_{a})$ plane. See also [20] for a full Bayesian analysis where it is found that this type of non-minimally coupled dark energy model is strongly favoured over $\Lambda$CDM. However, in this paper, we will focus on the phenomenology and whether or not the aforementioned dark energy parametrizations are equipped to capture it.
@@ 5.2 Reproducing dark energy phenomenology
As can be seen in Fig. (3), the fitting procedure we have used reproduces the predicted cosmological observables of both thawing quintessence and non-minimal thawing quintessence very accurately for all parameterizations considered. While there are some differences between the parameterizations, given that we have $\text{DoF}\simeq 20$ here, when examining the $\chi^{2}$ determined from minimizing Eq. (3.12) we consistently have that $\chi^{2}/\text{DoF}<1$ or $\simeq 1$ across the board regardless of the dark energy model or parameterization considered. For both the positive branch of thawing quintessence and the non-minimal thawing quintessence model, $\sim 99.99\%$ of the models for all parameterizations have $\chi^{2}/\text{DoF}<1$. For the hilltop branch it is a bit more interesting as the JBP parameterization is such that $\sim 99\%$ of the models have $\chi^{2}/\text{DoF}<1$, whereas for the other parameterizations it is $\sim 95\%$ of the models that have $\chi^{2}/\text{DoF}<1$.
@@ 5.3 Data constraints and viability of the dark energy models
This is confirmed when we compare the various parametrizations and thawing quintessence in terms of their ability to fit the complete (uncompressed) data. To do so, we find the best fitting model of thawing quintessence (i.e. that maximizes the likelihood ${\cal L}$ or, alternatively, minimizes $\chi^{2}=-2\ln{\hat{\cal L}}$ for the given datasets [^5] Here, we use the lite version of Planck likelihood to do the minimization as it has substantially fewer parameters than the full likelihood and is thus more efficient.). We do the same with $\Lambda$CDM and with the CPL, JBP, EXP, and BA parameterizations. The best-fit $\chi^{2}$ can be seen in Table (2) and show that both thawing quintessence and all parametrizations can fit the data better than $\Lambda$CDM, in agreement with [29]. In particular, while the scalar field improves the $\chi^{2}$ of $\Lambda$CDM only by $\Delta\chi^{2}=-2.9$, the parametrizations obtain $\Delta\chi^{2}=-6.7,\,-6.9,\,-7.1$, for CPL, EXP and BA. The worst performing parametrization is JBP, with $\Delta\chi^{2}=-4.9$, and this worse fit is the reason why in Fig. (4), for JBP, the wedge of the theory projection has a bigger overlap with the data constraints: both JBP and thawing quintessence are relatively close in terms of their ability to fit the expansion history data, meaning that when thawing quintessence is projected into JBP parameter space there will be more overlap than with the other parameterizations which are further away from thawing quintessence in terms of their ability to fit the data.
@@
There are, however, a couple of points worth briefly commenting on. Naively, if we use the Akaike Information Criterion (AIC), which is given by ${\rm AIC}=2k+\chi^{2}$, where $k$ is the number of parameters in the model, we find that $\Delta{\rm AIC}\simeq 12.5$ because the non-minimal model has 5 parameters (the 4 mentioned above plus one additional parameter $\dot{\phi}_{ini}$ for the initial conditions) while $\Lambda$CDM has only one ($\Lambda$). Even accounting for the additional number of parameters, this is still a notable improvement that, as mentioned above, merits further investigation. See [20] for further discussion and statistical analysis.
Another point to consider is that, while Section (5.2) demonstrated that the various dark energy parameterizations can all accurately reproduce the phenomenology of both minimal and non-minimal thawing quintessence, the fact that the non-minimal model can achieve a lower $\chi^{2}$ than any of the parameterizations does show that there is some information that is being lost when the full dark energy model is projected into the parameterizations’ parameter spaces. However, this makes sense as fitting directly to the observables using Eq. (3.12) was only fit to background observables that represent compressed versions of the full data (with the exception of the BAO data which requires no compression). The biggest difference between the compressed (background) observables and the full data will be in the CMB power spectra and CMB lensing data from Planck and ACT, where one must compute the full linear cosmological perturbations to compare with the data. And indeed, while we find that the non-minimal model reproduces similar $\chi^{2}$ with respect to the SNe and BAO components of the full data when compared with the various parameterizations, the improvement in overall $\chi^{2}$ ($\Delta\chi^{2}\simeq 4$ in the case of the DES-Y5 dataset) over those of the various parameterizations was largely driven by the Planck and ACT components of the data. So while these parameterizations clearly perform exceptionally well at reproducing dark energy phenomenology at the background level, it is perhaps not surprising that there is some information lost at the perturbative level.
@@ 6 Discussion
Our analysis reveals that assessing the viability of these dynamical dark energy models is remarkably robust to the choice of parameterization. By mapping the dynamical dark energy models into the $(w_{0},w_{a})$ parameter space for each parameterization, we demonstrated both that our conclusions regarding their viability (or how well they describe the data) are not overly subject to particular choices of parameterization and also that all of these various parameterizations effectively capture the phenomenology of the full dark energy models. This demonstrates that, at least for these very broad and simple kinds of microphysical dark energy models considered here, any of these choices of parameterization are sufficient to interpret the data and make assessments regarding the viability of such dark energy proposals.
@@ Appendix A Compressed data for theory priors
This is exactly analogous to what is done in the study of inflation models. We make observations that constrain the tensor-to-scalar ratio $r$ and the scalar spectral index $n_{s}$, and then use these constraints to determine the observationaly favored region of the $(r,n_{s})$ plane. To evaluate whether or not specific inflationary models are viable, we compute their predictions for $(r,n_{s})$ and compare these to those constraints (e.g. [148, 149, 150, 151, 152]). However, unlike with inflation, as we have seen there are a number of distinct parameterizations for dark energy and these parameters are a bit further from the actual observables. So, in order to actually determine how well a dark energy model maps onto these constraints, we must determine what a dark energy model would predict for $(w_{0},w_{a})$ by using a procedure that mimics what was done for determining where the data itself lies in terms of the parameters. In the dark energy literature, it is common to generate a dark energy model’s equation of state and “fit” this equation of state over some recent range of redshifts using one of the parameterizations directly as a proxy for determining where a model might lie on the $(w_{0},w_{a})$ plane. However, as discussed in [82, 13, 12], because we do not measure the equation of state directly, doing this can often lead to highly misleading results. In other words, we must generate the actual observables for a given dark energy model, and then determine what $(w_{0},w_{a})$ parameters best reproduce those observables given the known errors and uncertainties of the data. This will allow us to make an “apples to apples” comparison.